OpenStudy (idealist10):

A cup of boiling water is placed outside at 1:00 PM. One minute later the temperature of the water is 152 degrees Fahrenheit. After another minute its temperature is 112 degrees Fahrenheit. Find the outside temperature.

OpenStudy (idealist10):

@ganeshie8 @hartnn @wio @perl @dan815 @TuringTest @Compassionate

OpenStudy (idealist10):

@Directrix @CGGURUMANJUNATH

OpenStudy (anonymous):

If the outside temperature is \(T_{out}\), then we have some equation the form of: \[ T(t) = (T_{0}-T_{out})e^{kt}+T_{out} \]

OpenStudy (anonymous):

We can say \(T_0\) is the initial temperature... the temperature of boiling water.

OpenStudy (anonymous):

Which is \(212^\circ F\), so:\[ T(t) = (212^\circ F-T_{out})e^{kt}+T_{out} \]So now we have two variables here... \(k\) and \(T_{out}\). We also have two equations. \[ T(1) = 152^{\circ}F = (212^\circ F-T_{out})e^{k}+T_{out}\\ T(2) = 112^{\circ}F = (212^\circ F-T_{out})e^{2k}+T_{out} \]

OpenStudy (anonymous):

Subtracting \(T_{out}\) from both sides of both equations and dividing the second equation by the first equation gives: \[ \frac{T_{out}-112^{\circ}F}{T_{out}-152^\circ F} = e^{k} \]

OpenStudy (anonymous):

Substituting that into the first equation gives:\[ 152^{\circ}F = (212^\circ F-T_{out})\left(\frac{T_{out}-112^{\circ}F}{T_{out}-152^\circ F} \right)+T_{out} \]

OpenStudy (idealist10):

How did you get 212 degrees Fahrenheit?

OpenStudy (anonymous):

Boiling water

OpenStudy (idealist10):

Thanks!